OpenStudy (starfireandrobin):

Hello everyone!! How do you solve -5x-6/2 ??

OpenStudy (13mysterious):

@563blackghost

OpenStudy (moldybubblegum11):

i dont know this but i do know someone who might be able to help. @AloneS @Nnesha

OpenStudy (13mysterious):

is it -5x minus 2 or -5 times -6/2?

OpenStudy (starfireandrobin):

\[\frac{ -5x-6 }{ 2 }\]

OpenStudy (13mysterious):

Give me a second,

OpenStudy (3mar):

To solve such expression, you need to equal this to zero. Then\[\frac{ -x-6 }{ 2 }=0\] then \[-x-6=0\] \[x=-6\]

OpenStudy (starfireandrobin):

What happened to the 5??

OpenStudy (3mar):

but I think the expression is \[-5x-\frac{ 6 }{ 2 }\] not \[\frac{ -5x-6 }{ 2 }\]

OpenStudy (3mar):

Anyway, here is the answer

OpenStudy (starfireandrobin):

It is the equation I said it is.

OpenStudy (starfireandrobin):

Actually, the whole equation is \[\frac{ 1 }{ 3 }(\frac{ -5x-6 }{ 2 })-\frac{ 7 }{ 8 }\]

OpenStudy (skullpatrol):

Where is the equals sign?

OpenStudy (starfireandrobin):

There's no equals sign!! That's just the equation!!

OpenStudy (starfireandrobin):

I'm trying to find out what expressions are equivalent to my equation.

OpenStudy (3mar):

sorry \[\frac{ -5x-6 }{ 2 }=0\] \[-5x-6=0\] \[x=-6/5=-1.2\]

OpenStudy (3mar):

Thanks, StarfireAndRobin.

OpenStudy (starfireandrobin):

Why does it need to be equaled to 0?

OpenStudy (skullpatrol):

All I'm saying is an EQUATION must have an "="

OpenStudy (starfireandrobin):

Well..It doesn't. I don't know what to tell you.

OpenStudy (starfireandrobin):

Unless this has a different name and I'm calling it "equation" out of habit.

OpenStudy (skullpatrol):

That is called a variable "expression" :-)

OpenStudy (starfireandrobin):

@563blackghost

OpenStudy (starfireandrobin):

Ahh @skullpatrol thank you for clarifying.

OpenStudy (starfireandrobin):

So is that why @3mar was trying to make it equal to 0?

OpenStudy (skullpatrol):

Yup

OpenStudy (starfireandrobin):

So can you help me with solving it?? @skullpatrol

OpenStudy (skullpatrol):

Obviously the last choice is one of the right answers, right?

OpenStudy (starfireandrobin):

Um, possibly? I'm not too sure, as it's just the expression with the 1/3.

OpenStudy (starfireandrobin):

I'm under the assumption that what they mean by "equivalent to the expression" is what comes next. Like, when you start solving it, the expression will turn into another version of itself, right? So that's why I'm trying to solve it and see if by solving it, I get one of those choices.

OpenStudy (starfireandrobin):

I don't know if I explained myself well enough, sorry.

OpenStudy (skullpatrol):

That is correct.

OpenStudy (starfireandrobin):

Ah, great. So, what they're doing is simply multiplying 1/3 by -5x-6/2??

OpenStudy (starfireandrobin):

Since in the last choice, the 2 in the bottom of -5x-6/2 becomes a 6.

OpenStudy (starfireandrobin):

@skullpatrol ???

OpenStudy (skullpatrol):

Yup

OpenStudy (skullpatrol):

Now what?

OpenStudy (starfireandrobin):

I don't know, I'm trying to figure it out.

OpenStudy (starfireandrobin):

Again, it's affirmative that the last choice is one of the answers?? I don't want to fail my test.

OpenStudy (skullpatrol):

Yes, it is affirmative.

OpenStudy (starfireandrobin):

Alright, alright..So, I'm not too sure how to continue. Could you help me??

OpenStudy (skullpatrol):

Are you writing a test?

OpenStudy (starfireandrobin):

I'm taking a test on my online school and I don't want to fail.

OpenStudy (skullpatrol):

Nobody wants to fail, but you are cheating.

OpenStudy (starfireandrobin):

I'm not cheating...I'm seeking help. I DO NOT put an answer without knowing how to solve it, even if I have the answer right in front of me.

OpenStudy (starfireandrobin):

I do not cheat. It's just that my unit test for my online school is giving me material that they have not taught me, and this is something that isn't recognizable. That's why I need some help from you to learn it, if you know what to do.

OpenStudy (starfireandrobin):

I'm not cheating, I'm learning. And I ask of you, since nobody else is helping me, and the math chat is so busy being off topic of the subject to even ask for help there, for you to help me and learn how to identify what is equivalent to this expression, so sometime, I can do this on my own.

OpenStudy (starfireandrobin):

@skullpatrol Are you still there??

OpenStudy (skullpatrol):

ok, I'll try to guide you...let's start by using the last answer.

OpenStudy (starfireandrobin):

Thank you.

OpenStudy (skullpatrol):

Don't thank me yet :-)

OpenStudy (skullpatrol):

Do you know what a common denominator is?

OpenStudy (starfireandrobin):

A shared multiple?

OpenStudy (skullpatrol):

Yup, use that idea to find one please.

OpenStudy (starfireandrobin):

On -5x-6/6?

OpenStudy (starfireandrobin):

For 6 to become 8?

OpenStudy (skullpatrol):

What is the shared multiple of 6 and 8?

OpenStudy (starfireandrobin):

I have a question..If the bottom choice is what happens when you multiply 1/3(-5x-6/2), then why is the calculator giving me this?

OpenStudy (skullpatrol):

Because 6/6 = 1

OpenStudy (starfireandrobin):

Ahhh, alright. With that, 6 and 8 both share 24, don't they?

OpenStudy (skullpatrol):

Yes they do.

OpenStudy (starfireandrobin):

So times -5x-6/6 by 3??

OpenStudy (skullpatrol):

Why 3?

OpenStudy (starfireandrobin):

AHH I meant 6 xd

OpenStudy (starfireandrobin):

I was thinking of 8 * 3 = 24

OpenStudy (starfireandrobin):

My bad.

OpenStudy (starfireandrobin):

No wait...I'm so sorry, I just confused myself 100%. I'm very stupid, I'm sorry.

OpenStudy (starfireandrobin):

I'm so upset I don't know what to do.

OpenStudy (skullpatrol):

relax and think

OpenStudy (starfireandrobin):

I've been thinking!!! I've been sitting here for more than 30 minutes!! I went through all my questions already, and this is question 3. I've postponed this question for a while because I couldn't figure it out and I still don't know.

OpenStudy (skullpatrol):

Where are you stuck?

OpenStudy (starfireandrobin):

I'm honestly still stuck on 1/3(-5x-6/2). I wouldn't even know how to go on from there, although I know I have to multiply.

OpenStudy (starfireandrobin):

I'm just trying to go with it the best I can but I'm not understanding.

OpenStudy (skullpatrol):

this 1/3(-5x-6/2) is not right

OpenStudy (starfireandrobin):

 That's what it loookkssss liiikeeeee